If the frequency of an electromagnetic wave is 6.0 x 108 Hz, what is its wavelength?


1
Expert's answer
2021-05-14T09:41:11-0400

The frequency and length of the electromagnetic wave are related by the formula

"\lambda =\frac{c}{\nu}"

Where "c=3 \cdot 10^{8}m/s" - speed of light

Then we write

"\lambda =\frac{c}{\nu} =\frac{3 \cdot 10^{8}m/s}{6 \cdot 10^{8}1/s}=0.5 m"